During August, Leming Inc. worked on two jobs. Data relating to these two jobs follow:

Job 64

Job 65

Units in each order 50 80
Units sold 50 --
Materials requisitioned $3,560 $785
Direct labor hours 410 583
Direct labor cost $6,720 $9,328

Overhead is assigned on the basis of direct labor hours at a rate of $11. During August, Job 64 was completed and transferred to Finished Goods. Job 65 was the only unfinished job at the end of the month.

Required:
1. Calculate the per-unit cost of Job 64.
2. Compute the ending balance in the work-in-process account.
3. Prepare the Aug. 31 journal entries reflecting the completion and sale on account of Job 64. The selling price is 175% of cost. Refer to the chart of accounts for account titles. Round all journal entry amounts to the nearest dollar.

1. Calculate the per-unit cost of Job 64.

The per-unit cost of Job 64 is .

2. Compute the ending balance in the work-in-process account.

The ending balance in the work-in-process account is .

3. Prepare the Aug. 31 journal entries reflecting the completion and sale on account of Job 64. The selling price is 175% of cost. Refer to the Chart of Accounts for the exact wording of account titles. Round all journal entry amounts to the nearest dollar.

PAGE 1

GENERAL JOURNAL

DATE ACCOUNT POST. REF. DEBIT CREDIT

1

2

3

4

5

6
Answer #1

Answer:

1) Per unit cost of job 64 = (3,560+6,720+(410*11))/50= $295.8 per unit
2) Ending work in process account = (785+9,328+(583*11)) = $16,526
3) GENERAL JOURNAL
Date Account and explanation debit($) credit($)
Aug. 31 Finished goods($295.8*50)            14,790
          Work in process          14,790
(To record completion of Job 64)
Aug. 31 Account Receivable (14,790*175%)       25,882.50
            Sales revenue     25,882.50
(To record sales of Job 64)
Aug. 31 Cost of goods sold            14,790
              Finished goods          14,790
(To record cost of goods sold for Job 64)